Stephanie wishes to make 6 litres of drink containing 60% orange juice by mixing plain water and concentrate which is 90% orange juice. How many litres of plain water and concentrate should she mix?
Answer by nerdybill(7384)

Stephanie wishes to make 6 litres of drink containing 60% orange juice by mixing plain water and concentrate which is 90% orange juice. How many litres of plain water and concentrate should she mix?
.
Let x = amount (litres) of plain water
then
6-x = amount (litres) of concentrate
.
.90(6-x) = .60(6)
5.4-.90x = 3.6
-.90x = -1.8
x = -1.8/(-.90)
x = 2 litres (water)
.
Concentrate:
6-x = 6-2 = 4 litres